OpenStudy (anonymous):

find the slope of tangentg line at x(1/x2 - 1/4) when x=-2

OpenStudy (anonymous):

is your function\[f(x) = \frac{ 1 }{ x }-\frac{ x }{ 4 }\]?

OpenStudy (anonymous):

When I distribute the x through the first equation I get: y=1/x - x/4 so y'= 1/x^-2 - 1/4 Plugging in x= -2 I get y'(-2) = -1/2 (this is the slope of the tangent line) In order to find the equation of the tangent line Plugging in x= -2 in the first equation I get y = 0 Using the Point-Slope formula: y-0 = -1/2(x-(-2)) So the equation of the tangent line is y = -1/2x - 1

OpenStudy (anonymous):

your y' is off by a little bit. you have y = x^(-1) - x/4 y' = -x^(-2) - 1/4 Try recalculating your slope based on this

OpenStudy (anonymous):

@jollyjolly0 thanks @jonathan1992v y' is infact -x^(-2) - 1/4 but I just forgot include the (-) sign, but in my calculations I included it, so the slope stays the same.